Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The PJX5 will reduce costs by squeezing more juice from each plum and doing so in a more efficient manner. Mr. Bensen gave Derek the following information. What is the NPV of the PJX5?

a. The PJX5 will cost $1.71 million fully installed and has a 10 year life. It will be depreciated to a book value of $285,571.00 and sold for that amount in year 10.

b. The Engineering Department spent $23,384.00 researching the various juicers.

c. Portions of the plant floor have been redesigned to accommodate the juicer at a cost of $17,970.00.

d. The PJX5 will reduce operating costs by $369,007.00 per year.

e. CSD’s marginal tax rate is 33.00%.

f. CSD is 58.00% equity-financed.

g. CSD’s 18.00-year, semi-annual pay, 5.41% coupon bond sells for $969.00.

h. CSD’s stock currently has a market value of $20.82 and Mr. Bensen believes the market estimates that dividends will grow at 2.70% forever. Next year’s dividend is projected to be $1.60.Caspian Sea Drinks is considering the purchase of a plum juicer – the PJX5. There is no planned increase in production. The P
Answer #1

Solution:-

Depreciation Per year = \frac{1,710,000 - 285,571}{10}

Depreciation Per year = $142,442.9

To Calculate WACC-

Cost of Equity = \frac{Dividend}{Price} + Growth

Cost of Equity = \frac{1.60}{20.82} + 0.027

Cost of Equity = 10.38%

Weight of Equity = 58%

Weight of Debt = 1-0.58 = 42%

Cost of Debt-

Annual Cost of debt after tax = 2.84% * 2 * (1-0.33)

Annual Cost of debt after tax = 3.81%

WACC = 10.38% * 0.58 + 3.81% * 0.42

WACC = 7.62%

Note:- Engineering Department Cost and Redesigned Cost are Irrelevant Cost.

To Calculate NPV of the Project-

B с D E F G H J K Statement of Cash flows After Tax (Formula Sheet) 4 15 6 1 12 3 18 9 10 0 =-1710000 А 1 Solution :- 2 3 4 Y

NPV of the Project is $435,693

If you have any query related to question then feel free to ask me in a comment.Thanks.